TCS Testing + LDG Commissioning: Performance Verification - 22 April 2010 UT

Night 4:

Observer: DMiller (Tucson) JHill (LBTO)
Telescope Operator: JHill (LBTO)
Telescope Support: JLittle, BCurtis
Instrument Support: JMorris

Summary

Lost to weather: wind, humidity, clouds, rime, snow (your choice)

Preparations

AGW2 is reconnected and operational again (although we aren't planning to use it tonight).

JHoward adjusting elevation pad pressures in the afternoon.

Butch switched DX Primary over to the old compressor. This has the side effect of better behaviour from the aftercoolers supplying only one mirror.

Details

01:40 JMorris fills LBCs.

03:05 Changed HBS LoopTempOffset from my special value of -12 degC to -9 degC to be sure that the oil isn't overcooled while James is adjusting pad pressures. (Nominal value is -10 degC.)

04:00 James and John adjust the "K" factor in HBS from 170 to 125 in order to bring the system pressure back down to 120 bar. This should noticeably improve the hot oil situation, but could make friction worse.

04:30 Wind gusts above 30 m/sec. Humidity 102%. T=-4.5 degC. (a lovely spring evening on Mt. Graham)

04:48 Restart MCSPU because of some wierdness in the AZ/EL DSP state. ENV thought enclosure azimuth was 1120 deg, and there were strange question marks in the MCSPU display.

05:50 THahn is running AGW software tests from Potsdam. So we are making some use of this snowy night.

06:01 Turned on shutter door and drain pipe heat tracing. Outside temperature is down to -6 degC.

-- DougMiller - 19 Apr 2010
Topic revision: r4 - 22 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback